The product of two numbers is 330. The smaller number is seven less than the larger number. What are the numbers?
Answer by ankor@dixie-net.com(22740)

The product of two numbers is 330. The smaller number is seven less than the
larger number. What are the numbers?
:
Two numbers x, y
:
"The product of two numbers is 330."
x*y = 330
:
"The smaller number is seven less than the larger number."
y = x-7
;
What are the numbers?
:
Substitute (x-7) for y in the 1st equation, solve for x
x(x-7) = 330
x^2 - 7x = 330
x^2 - 7x - 330 = 0
Factor
(x-22)(x+15) = 0
Two solutions
x = +22 then y = +15
x = -15 then y = -22
:
You can check them in the original equations